There are no Open Network Issues Currently

Server Status

Below is a real-time overview of our servers where you can check if there's any known issues.

Server Name HTTP FTP POP3 PHP Info Server Load Uptime
web2.hippynet.co.uk PHP Info
stream2 migration to MediaCP (Resolved) Critical

Affecting Server - stream2.hippynet.co.uk

  • 23/03/2021 02:00 - 18/04/2021 20:47
  • Last Updated 17/03/2021 11:42

We will be migrating stream2.hippynet.co.uk from Centovacast to MediaCP at 02:00 Tuesday 23rd march 2021. During the migration there will be a short disruption. StreamMachine customers will automatically fall back to their backup server. Although we do not anticipate any problems after migration we ask all customers to please check their streams and open a support ticket under high priority if there are any problems.

Emergency maintenance - Liverpool (Resolved) Critical

Affecting Other - All Liverpool services

  • 10/06/2019 21:07 - 11/06/2019 08:45
  • Last Updated 10/06/2019 21:08

Dear Customer,

This communication is to notify you of the following planned works which will begin shortly;

Service Affected: Liverpool

As previously advised

Start of maintenance:

10/06/2019 - 22:00 BST

End of maintenance:

11/06/2019 - 06:00 BST

Duration of impact:

Up to 3 hours disruption 

Impact of work:

Services will be hard down for 15 minute sessions

Classification of works:

Standard.

Description of works:

Following a recent Major Incident impacting a core router in Manchester, these works are to perform a hardware change and perform a vendor recommended operating system update to ensure future stability.

Should you have any questions on this notification please contact our NOC on our contact details listed below. In order to assist us in quickly finding your details please quote our reference number in any communications.

 Thanks

SSL warning on Stream2 (Resolved) Medium

Affecting Server - stream2.hippynet.co.uk

  • 29/05/2019 19:51 - 05/06/2019 16:02
  • Last Updated 05/06/2019 16:02

29/05/2019 - SSL Warning on stream2.hippynet.co.uk

Issue: expired SSL certificate on the Centovacast control panel

Impact on broadcast: None

The SSL certificate has expired on stream2.hippynet.co.uk and we are waiting on our certificate authority to issue a new SSL certificate.

You will get a warning when trying to access the Centovacast control panel on stream2.hippynet.co.uk. It is perfectly safe to click on continue and access the control panel ignoring the warning.

This issue will be resolved once the certificate authority issues a new SSL certificate.

05/06/19 - A new SSL certificate has been issued and this issue is now resolved.

Connectivity Issues from some locations (Resolved) Medium

Affecting System - Liverpool Cloud

  • 14/03/2018 23:18 - 11/05/2018 11:02
  • Last Updated 14/03/2018 23:25

14/03/2018 @ 22:00 - We are aware of connectivity issues from some locations. This issue is outside of our network and we are awaiting a resolution.

Incoming bandwidth to Liverpool Cloud (Resolved) High

Affecting System - Liverpool Cloud

  • 09/03/2018 00:10 - 09/03/2018 22:37
  • Last Updated 09/03/2018 00:11

08/03/2018 - 23:34 - We are currently seeing an issue with inbound bandwidth to the Liverpool datacentre. Live streaming customers and customers uploading media may see issues.

EU Packet loss (Resolved) High

Affecting System - Liverpool Data Centre

  • 23/04/2017 09:00 - 23/04/2017 15:56
  • Last Updated 23/04/2017 09:39

23/04/17 @ 09:00 - We are seeing packet loss on our European ring. Engineers are investigating.

stream2.hippynet.co.uk scheduled upgrade (Resolved) Medium

Affecting Server - stream2.hippynet.co.uk

  • 05/05/2017 02:00 - 05/05/2017 03:11
  • Last Updated 20/04/2017 04:26

We have scheduled downtime that will affect stream2.hippynet.co.uk. Stream from this server will be unavailable during this time. Customers with our Stream Machine add on will still stay on air as the backup stream will automatically take over.

Reason for outtage: To upgrade Cenovacast from v2 to v3
Scheduled window: Friday 5th May 02:00 - 06:00
Expected time to complete: 2 hour

If you have any questions please open a support ticket and we will be more than happy to help.

The Hippynet Team

Packet loss from Liverpool Data Centre (Resolved) High

Affecting Other - Liverpool Cloud

  • 10/01/2017 15:44 - 02/02/2017 21:54
  • Last Updated 10/01/2017 23:49

We are currently seeing packet loss from our Liverpool Data Centre. If you are hearing audio drop outs then this is the most likely cause. We are working to fix this as quickly as possible.

update 23:47 10/01/17
Traffic has been temporarily rerouted. Packet loss has been resolved. A permanent reolution will be made tomorrow.

Centovacast upgrade issues (Resolved) Critical

Affecting Server - stream2.hippynet.co.uk

  • 19/05/2014 11:13 - 20/05/2014 17:50
  • Last Updated 19/05/2014 11:15

11:15 GMT - We are seeing some problems caused by the upgrade to Centovacast V3. We are rolling back to a backup of V2 and awaiting further information from Centova.

Listener stats errors (Resolved) Medium

Affecting Server - stream1.hippynet.co.uk

  • 19/03/2014 16:23 - 27/03/2014 21:24
  • Last Updated 19/03/2014 16:24

Centovacast on Stream 1 is sometimes showing audio source disconnected and no listeneres present. This is being investigated.

Connectivity issues - Liverpool (Resolved) High

Affecting Other - Network - Liverpool

  • 04/08/2013 13:11 - 09/08/2013 13:15
  • Last Updated 04/08/2013 13:13

13:00 - We are currently seeing network connectivity issues in our Liverpool data centre. Engineers are working to correct this. Servers affected are:
web2.hippynet.co.uk
stream2.hippynet.co.uk
ns7.hippynet.co.uk
ns8.hippynet.co.uk

Scheduled tasks (Resolved) Medium

Affecting Server - stream2.hippynet.co.uk

  • 23/05/2013 14:06 - 23/05/2013 14:31
  • Last Updated 23/05/2013 14:14

Scheduled jobs are not running on Stream2. This affects statistics and server monitoring. The server will be rebooted as part of the process to rectify this problem.

The server has restarted and the scheduling problem has been rectified.

Server performance issues (Resolved) Medium

Affecting Server - stream2.hippynet.co.uk

  • 17/03/2013 23:36 - 02/04/2013 10:50
  • Last Updated 18/03/2013 00:59

We have identified a software bug in the stream source software running on Stream 2. Administrators are working on a fix for the bug.

  • Server reboots may be required.
  • Some customers may receive emails notifying them that their servers have been restarted.

Centovacast control panel access (Resolved) Critical

Affecting Server - stream2.hippynet.co.uk

  • 11/03/2013 00:33 - 11/03/2013 05:49
  • Last Updated 11/03/2013 05:49

Customers are currently experiencing issues accessing the Centovacast control panel.


00:37 - A software update has caused problems with access to Centovacast control panel. Our administrators are working to fix this.

01:48 - Access to Centovacast has been restored.

Powered by WHMCompleteSolution